Flounder Inc. uses a calendar year for financial reporting. The company is authorized to issue 9,680,000 shares of $10 par common stock. At no time has Flounder issued any potentially dilutive securities. Listed below is a summary of Flounder’s common stock activities.

1.

Number of common shares issued and outstanding at December 31, 2018

2,000,000

2.

Shares issued as a result of a 10% stock dividend on September 30, 2019

200,000

3.

Shares issued for cash on March 31, 2020

1,940,000

Number of common shares issued and outstanding at December 31, 2020

4,140,000

4.

A 2-for-1 stock split of Flounder’s common stock took place on March 31, 2021

  1. Compute the weighted-average number of common shares used in computing earnings per common share for 2019 on the 2020 comparative income statement.

  1. Compute the weighted-average number of common shares used in computing earnings per common share for 2020 on the 2020 comparative income statement.

  1. Compute the weighted-average number of common shares to be used in computing earnings per common share for 2020 on the 2021 comparative income statement.

  1. Compute the weighted-average number of common shares to be used in computing earnings per common share for 2021 on the 2021 comparative income statement.
Answer #1
a] As stock dividends have to be given effect to from the beginning of the year, the weighted average # shares = 2,000,000+10%*2,000,000 = 2200000
b] Weighted average number of shares = 2200000*3/12+4140000*9/12 = 3655000
c] Weighted average number of shares = 3655000*2 = 7310000
d] Weighted average number of shares = 4140000*2 = 8280000